OpenStudy (anonymous):

Which of the following options results in a graph that shows exponential growth? f(x) = 0.4(3)x f(x) = 3(0.5)x f(x) = 0.8(0.9)x f(x) = 0.9(5)−x

OpenStudy (anonymous):

Which of the following options results in a graph that shows exponential growth? f(x) = 0.4(3)x f(x) = 3(0.5)x f(x) = 0.8(0.9)x f(x) = 0.9(5)−x @Astrophysics @amoodarya

OpenStudy (astrophysics):

can you take an image

OpenStudy (anonymous):

i dont how

OpenStudy (astrophysics):

\[f(x) = 0.4(3)x ~~~~f(x) = 0.4(3)^x\]

OpenStudy (anonymous):

i using f(x)

OpenStudy (astrophysics):

which one is it?

OpenStudy (anonymous):

f(x)

OpenStudy (astrophysics):

What?

OpenStudy (anonymous):

f(x) = 0.4(3)

OpenStudy (astrophysics):

No...is the x an exponent or not

OpenStudy (anonymous):

yes

OpenStudy (astrophysics):

 is this how it is?